z^4-9=0 I need four answers!
Found 2 solutions by Alan3354, Edwin McCravy:
Answer by Alan3354(69443)

z^4-9=0
---------
z^4 = 9
z^2 = 3
Ans 1 z = sqrt(3)
Ans 2 z = -sqrt(3)
-------------------
z^2 = -3
Ans 3 z = i*sqrt(3)
Ans 4 z = -i*sqrt(3)

Answer by Edwin McCravy(20055)

            z4 - 9 = 0  

        (z2)2 - 32 = 0

  (z2 - 3)(z2 + 3) = 0

z2 - 3 = 0;   z2 + 3 = 0
    z2 = 3;       z2 = -3
     z = ±√3;      z = ±√-3
                   z = ±i√3

Four answers √3, -√3, i√3, -i√3

Edwin